The following equation shows the pork demand in China where I represent income and PC represents the price of chicken.

Qd = 500-4P+PC+0.3I (Q is in 1000 pound)

a) Find the demand equation when PC = 900 and I= 12000.

b) If P=800, what is the price elasticity of demand for the pork? (2)

c) Given 13.2 and 13.3, what is the income elasticity of demand?

d) Given 13.2 and 13.3, What is the cross-price elasticity of demand for pork and chicken?

e) Are pork and chicken demand substitute or demand complements? Would sales of pork be increase or decrease if price of chicken falls?
Answer #1

I 500- 4P + e + 0.3 Oo - 900 (1) 12000 O - - o = 500-4P + 1900) + 0.3 (12000) Soo- up + 4500 5000-YP & benes Demand equation(c) Od = 500 - 4P + P + 0.37 In come elasticity of demand lome By I we get ? putting value of p, f, and Od = 1800Income elasticity 2ooo (d) Cross-elasticity of demand for Pork and chicken 0 = 500*-up + P + 0. 31 by putting all values of Pis rel The cross- elasticity of demand Il for pork and chicken positive this implies pork and chicken are substitutes. So As